OpenStudy (anonymous):

An arrow is shot by an archer at a 50.0 degree angle at a velocity of 10.0 m/s. What is the maximum height and range of the arrow? **how do you calculate this? thank you:)

OpenStudy (tylerd):

we want the hieght when the final velocity in the y direction is 0

OpenStudy (anonymous):

yes:)

OpenStudy (tylerd):

so Vf=Vi+at

OpenStudy (tylerd):

but we are dealing with the y direction so

OpenStudy (anonymous):

okay:)

OpenStudy (tylerd):

10.0sin(50) = initial velocity in the Y direction

OpenStudy (tylerd):

the acceleration is from gravity so its negative which is a constant -9.8m/s^2

OpenStudy (tylerd):

well actualy i think we can use a different equation for htis to make it easier

OpenStudy (anonymous):

okie:)

OpenStudy (tylerd):

Vf^2=Vi^2+2ay i think it is

OpenStudy (tylerd):

0=(10sin(50))^2-2(9.8)(Y)

OpenStudy (tylerd):

solve for Y

OpenStudy (tylerd):

-[10sin(50)]^2=-2(9.8)(Y)

OpenStudy (tylerd):

10sin(50)/(2)(9.8) = Y

OpenStudy (anonymous):

okay, so Y=-12.856?

OpenStudy (tylerd):

lul wait wait

OpenStudy (anonymous):

haha okie

OpenStudy (tylerd):

oh i see

OpenStudy (tylerd):

i forgot to square it one sec

OpenStudy (anonymous):

ok!

OpenStudy (tylerd):

im getting 3m

OpenStudy (anonymous):

ohh okay! so this is the max height correct? 2.99=3 is the maximum height?

OpenStudy (tylerd):

ya

OpenStudy (anonymous):

awesome!! how do we find the range now?

OpenStudy (tylerd):

u mean where the arrow lands?

OpenStudy (tylerd):

in this case we need to find time.

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

how do we do that?

OpenStudy (tylerd):

so 10cos(50) = velocity in the x directiion

OpenStudy (tylerd):

ah i see so we need to find when the ball hits the ground

OpenStudy (tylerd):

to do that we have to consider whats happening in the Y direction first

OpenStudy (anonymous):

okie:)

OpenStudy (tylerd):

we already solved its max height

OpenStudy (tylerd):

now the time it hits the ground is equal to 2 times the time it takes for it to reach the max height

OpenStudy (anonymous):

yes:) 3.00!

OpenStudy (tylerd):

so we use the kinematic equation y=Vi(t)+1/2at^2

OpenStudy (tylerd):

or actually better to use

OpenStudy (tylerd):

Vf=Vi+at

OpenStudy (anonymous):

okay! so what do we plug in?

OpenStudy (tylerd):

so 0=10sin(50)-9.8(t)

OpenStudy (tylerd):

-10sin(50)=-9.8t

OpenStudy (tylerd):

10sin(50)/9.8 = t

OpenStudy (tylerd):

however this t corresponds to the time when the arrow is at its max height

OpenStudy (tylerd):

and we want the time that is twice that

OpenStudy (anonymous):

ahh okay, and the t=-0.2677?

OpenStudy (tylerd):

t = 1.56

OpenStudy (tylerd):

well the t in the equation is half that

OpenStudy (tylerd):

but we multiply it by 2 to get the time for it to hit the ground rather then the time to reach its max height

OpenStudy (tylerd):

so then we have x=Vi(t)

OpenStudy (anonymous):

ohh okay!

OpenStudy (tylerd):

x=10cos(50)(1.56)

OpenStudy (tylerd):

should give u ure answer

OpenStudy (anonymous):

ahh okie:) so x=15.053 that is the range?

OpenStudy (tylerd):

x=10 10cos(50)(1.56)

OpenStudy (anonymous):

ohh okay! yay!! so mx height=3 and range=10 ?

OpenStudy (tylerd):

ya

OpenStudy (anonymous):

yay! thank you !! :)